OpenStudy (anonymous):

An acute angle θ is in a right triangle with sin θ = 3/4 . What is the value of cot θ?

OpenStudy (raden):

 use pythagorean theorema to find of x next, just calculate : cot(theta) = x/3 = ... / 3

OpenStudy (anonymous):

3/scientific notation 7

OpenStudy (raden):

nopes... x^2 = 4^2 - 3^2 x^2 = 16 - 9 x^2 = 7 x = sqrt(7)

OpenStudy (anonymous):

sqrt(7)/3

OpenStudy (raden):

yes

OpenStudy (raden):

for cot(theta)

OpenStudy (anonymous):

thanks again

OpenStudy (raden):

welcome